Fermi National Laboratory


Accelerator Update - August 8-August 11

Friday August 8
The day shift started with AntiProton Source (Pbar) stacking, beam to Miniboone, and Tevatron (Tev) studies.

At 10:00 AM problems with MECAR kept beam off to the Main Injector (MI) and the Tev for ~ 40 minutes.

At 2:54 PM Tev studies were halted by a quench at A28U, the quench was due to an error writing tables to AQ7.

Shot setup for the Tev began at 8:00 PM, with store #2879 colliding by 10:42 PM with an initial average luminosity of 40.50e30.

Saturday August 9
The midnight shift started with store #2879 colliding, stacking and beam to Miniboone.

At 5:00 AM the store was aborted due to a Tev RF anode inhibit. The cause was determined to be the Longitudinal Damper System driving the RF in an unpleasant way. The system was physically disconnected at F0.

Tev studies started at 6:30 AM.

Shot setup for the Tev began at 12:29 PM, with store #2883 colliding by 2:15 PM with an initial average luminosity of 42.15e30.

We then returned to stacking.

Tev losses were very high, approximately one hour of standard tuning was required.

Switchyard, Slipstacking, and Numi studies began approximately at 11:00 PM.

Sunday August 10
The midnight shift started with store #2883 colliding, stacking, beam to Miniboone and various studies.

Switchyard 120 studies concluded at 1:00 AM.

Slipstacking studies finished at 3:00 AM.

Numi studies were completed by 4:30 AM.

Store #2883 was lost at 5:06 AM, when F48L quenched.

Quench recovery was complete by 7:30 AM. Cause of the quench appeared to be a bad ramp card.

Shot setup began at 12:03 PM.

Shot setup was delayed due to high emmitances in Pbar.

Store #2887 was colliding by 4:43 PM with an initial average luminosity of 48.88e30. This is a new record.

Once again Tev losses were high and some tuning was required.

Pbar returned to stacking.

Monday August 11
Store #2887 was lost at 12:05 AM due to a F4 lead failure.

Tev was back up by 3:00 AM. Some changes were made to the control loops for F4.

Shot setup began at 7:24 AM.

Today's Plan
Put store in Tev, stacking and sending beam to MiniBoone.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/11/2003   email Fermilab